Friday, December 5: OWL Shift

From Hall A Wiki
Jump to: navigation, search

Go back to Daily Run Plan Daily Run Plans

  • General considerations
    • Tanja Horn is the current Run Coordinator. Please contact directly via RC phone 757-876-1787
    • Follow the generic Instructions for shift takers
    • Shift leaders: Please make sure to fill out the beam time accounting form and write a "Shift Summary" log entry at the end of your shift regardless if there is beam or not.
    • In general, Hall technical staff should not be contacted by the shift crew - please call the RC first.
  • Instructions for this shift
    • The target operator is required to supervise the cryo-target.
    • Kinematics, DVCS Kin1: Ebeam = 7.3 GeV, PLHRS=2.71 GeV, Theta_HRS=22.83, Theta_calo=12.35
    • Make sure to use the DVCS CODA configuration for all runs (even when taking single arm HRS data).
    • Continue the following measurements:
      • 5 uA unrastered beam at the nominal center of the carbon target
      • Turn off auto-gains on Hall BPMs and take a run with variable beam currents between 2-5 uA (0-5-0-2-0) to check the BPM pedestals.
      • BPM Calibrations (Bullseye Scan) - do this even if scans from the two harps are not available. Instructions of the procedure can be found here.
      • Take a run with the optics target with unrastered beam at 5 uA [~30 minutes]
      • Take runs with LH2 target at 5 uA, 2x2 raster. Keep taking these runs.
    • Note that there may be a cryo switch from ESR to CHL at some point late in the shift which requires the magnets to be ramped down. see this logbook entry for details on the cryo sitation